# Quantum Long Short-Term Memory (QLSTM) vs Classical LSTM in Time Series Forecasting: A Comparative Study in Solar Power Forecasting

### v2

@@
Abstract
Abstract: Accurate solar power forecasting is pivotal for the global transition towards sustainable energy systems. This study conducts a meticulous comparison between Quantum Long Short-Term Memory (QLSTM) and classical Long Short-Term Memory (LSTM) models for solar power production forecasting. The primary objective is to evaluate the potential advantages of QLSTMs, leveraging their exponential representational capabilities, in capturing the intricate spatiotemporal patterns inherent in renewable energy data. Through controlled experiments on real-world photovoltaic datasets, our findings reveal promising improvements offered by QLSTMs, including accelerated training convergence and substantially reduced test loss within the initial epoch compared to classical LSTMs. These empirical results demonstrate QLSTM’s potential to swiftly assimilate complex time series relationships, enabled by quantum phenomena like superposition. However, realizing QLSTM’s full capabilities necessitates further research into model validation across diverse conditions, systematic hyperparameter optimization, hardware noise resilience, and applications to correlated renewable forecasting problems. With continued progress, quantum machine learning can offer a paradigm shift in renewable energy time series prediction, potentially ushering in an era of unprecedented accuracy and reliability in solar power forecasting worldwide. This pioneering work provides initial evidence substantiating quantum advantages over classical LSTM models while acknowledging present limitations. Through rigorous benchmarking grounded in real-world data, our study illustrates a promising trajectory for quantum learning in renewable forecasting.
keywords:
Quantum Machine Learning, Forecasting, Renewable energy systems, Quantum Neural Networks
@@ 1 Introduction
Despite these promising developments, the existing literature significantly lacks a comprehensive, empirical comparison between QLSTM and classical LSTM models grounded in real-world solar production data. Preceding works 10.5555/1659652.1659656 9892441 have primarily focused on synthetic benchmarks or theoretical analysis of Quantum Recurrent Neural Networks variants, leaving a critical gap in our understanding of the practical implications of QLSTMs for renewable energy forecasting. This investigation aims to bridge this divide, offering a thorough comparative analysis in this pivotal domain.
@@ 2 Background or Related Work
The global energy landscape is undergoing a transformative shift towards sustainable and renewable solutions, driven by the pressing need to mitigate the environmental impact of traditional fossil fuel-based systems. Solar power has emerged as a pivotal force in reshaping energy production and consumption patterns, offering a promising pathway to mitigate the environmental impact of traditional fossil fuel-based systems. However, the large-scale integration of intermittent renewable sources into existing infrastructure poses multifaceted challenges that must be addressed to facilitate a seamless and efficient transition. Inaccuracies in forecasting the generation of solar and wind power can lead to significant deviations from planned electricity schedules, resulting in imbalances, inefficiencies, and substantial costs for grid operators and utilities GOODARZI2019110827 . Consequently, policymakers have prioritized improved renewable forecasting to mitigate such challenges.
@@
Amidst this paradigm shift, the convergence of quantum information (QI) and machine learning (ML) has ushered in a revolutionary approach to data analytic: Quantum Machine Learning (QML) dunjko2017machine . This paradigm-shifting synthesis harnesses techniques from both quantum computing and traditional machine learning, offering innovative solutions to longstanding obstacles across diverse sectors, including renewable energy Biamonte2017 . Significantly, QML transcends mere energy minimization tasks, presenting a broader scope in problem-solving paradigms NEURIPS2020_0ec96be3 , thereby unlocking new avenues for precise solar power predictions.
@@
As the large-scale penetration of renewable sources necessitates proactive management of electrical grids, advanced prediction methodologies for intermittent energy sources, particularly photovoltaic plants, have become paramount succetti2020 ; prema2015 . Accurate solar power forecasting plays a critical role in enabling grid operators to maintain a delicate equilibrium between energy creation and utilization. Notably, solar production forecasting over longer time horizons does not mandate real-time predictions, providing an opportunity where the potentially slower inference times of quantum models may be acceptable in exchange for substantially improved accuracy.
@@
In recent years, there has been a surge of interest in quantum machine learning developing and refining of quantum adaptations of recurrent neural networks (RNNs) for time series forecasting applications 10.1007/978-3-031-19493-1_6 lindsay2023 , thereby substantiating the potential of quantum computational models in predictive analytics. Marking a pivotal shift, the seminal work of Chen et al. 9747369 pioneered the introduction of QLSTM architectures, which amalgamate variational quantum circuits Cerezo_2021 with the conventional LSTM framework. This innovative synthesis harnesses an exponentially larger Hilbert space for data representation and computation, potentially enabling the capture of higher-order correlations and intricate temporal dynamics.
@@
While classical Long Short-Term Memory (LSTM) neural networks have demonstrated remarkable efficacy in leveraging long-term temporal dependencies for accurate forecastingZHENG2020114001 sorkun2020 succetti2020 meenal2022 , they struggle to capture the complex, non-linear spatiotemporal patterns inherent in solar power generation LINDEMANN2021650 , which involve intricate relationships between meteorological variables, solar irradiance, and power output. Quantum Long Short-Term Memory (QLSTM) networks, which leverage the principles of quantum mechanics and an exponentially larger Hilbert space, hold the potential to address these limitations by enabling more effective mapping of these intricate relationships between weather variables, solar irradiance, and power generation more effectively. and capturing higher-order correlations and temporal dynamics.
@@ 3 Research Motivation and Novelty
This advancement, coupled with the promise of faster convergence times and heightened resilience to noise emmanoulopoulos2022quantum , paves the way for the realization of more precise and reliable forecasting systems – a quality of paramount importance in the domain of solar power production. However, it is crucial to acknowledge that the nascent field of quantum machine learning presents its own unique challenges and opportunities Cerezo2022 . These challenges encompass the current hardware limitations, the potential trade-offs between quantum advantages and computational overhead, as well as the need for systematic optimization and validation across diverse scenarios – factors that this study meticulously considers in its comparative analysis.
@@ 4 Contributions
  • 1.
    Empirical Validation on Real-World Data: Providing the first comprehensive empirical evidence that validates the utility of Quantum Long Short-Term Memory (QLSTM) networks for solar power forecasting. This study transitions from synthetic benchmarks to operational photovoltaic plant datasets, offering a realistic evaluation of quantum architectures on genuine renewable time series data. Quantitative results demonstrate that QLSTMs achieve up to Ximprovement in accuracy and Yfaster convergence compared to their classical LSTM counterparts.
  • 2.
@@
  • 5.
    Establishing a Credible Alternative: By benchmarking against classical techniques and conventional neural networks using actual photovoltaic data, this investigation establishes QLSTMs as a credible alternative to traditional forecasting methods. The results accentuate the potential of quantum machine learning in fortifying the accuracy and reliability essential for sustainable energy infrastructure planning and execution, enabling improvements of up to Zin forecasting accuracy compared to state-of-the-art methods.
  • 6.
@@ 5.1 Data Description
Table 1: Comparison of Real-World and Simulated Solar Power Data.
@@ 5.3 Simulation Framework
Our exploration of Quantum Long Short-Term Memory (QLSTM) models was made possible using the PennyLane quantum machine learning framework bergholm2022pennylane . PennyLane, at its core, blends quantum and traditional computing to help build and refine models, benefiting from its ability to automatically adjust model parameters.
@@
Bridging the gap between the quantum and regular sections, PennyLane’s qml.qnn.TorchLayer connects the quantum elements with the regular PyTorch framework NEURIPS2019_bdbca288 . This ensures a smooth flow of adjustments during the optimization phase.
@@ 5.4 Architecture
The LSTM and QLSTM architectures are detailed in Appendix A. This research encompasses the design and deployment of both LSTM and QLSTM architectures. These architectures were judiciously crafted to allow a fair comparison between classical and quantum techniques, with a focal point on solar power forecasting.
@@
Diverging, the QLSTM model replaces the classical LSTM cells with variational quantum circuits (VQCs), an implementation adapted and enhanced from qlstm repository for parts of speech tagging rdisipio . This substitution aims to harness the computational advantages unique to quantum mechanisms. Echoing the LSTM’s design, the QLSTM layers two of these quantum circuits. The VQCs, in their capacity as quantum feature extractors, exploit the deep representational capabilities of quantum states, encoding intricate time series dynamics. These parametric circuits oscillate between rotation and entanglement gates, ensuring a concise representation of temporal patterns within the exponentially expansive Hilbert space. The qubit quantity and circuit depth were adapted in alignment with the specific characteristics intrinsic to the solar forecasting data. Notably, outside of its quantum encoding, the QLSTM’s broader workflow aligns seamlessly with the LSTM’s, facilitating a direct comparison. A dropout rate of 0.2 is also infused between the QLSTM layers, maintaining consistency.
@@
  • 2.
    Quantum Variational Circuit (VQC): Data is encoded into the quantum circuit using rotation operations, followed by alternating entanglement and variational rotation layers detailed in Appendix A.
  • 3.
@@
1
Input layer
Input layer
2
LSTM layer (with input, output, forget gates and memory cell)
QLSTM layer (with VQCs featuring rotation and entanglement gates)
3
Dropout (0.2 rate)
Dropout (0.2 rate)
4
LSTM layer (with input, output, forget gates and memory cell)
QLSTM layer (with VQCs featuring rotation and entanglement gates)
5
Dropout (0.2 rate)
Dropout (0.2 rate)
6
Linear output layer
Linear output layer
Table 2: Comparison of LSTM and QLSTM Architectures.
@@ 6.1 Statistical Analysis
The statistical analysis of the train loss, depicted in Table 3, suggests a potential trend where the QLSTM model gradually outperforms its classical counterpart. Despite the p-value slightly exceeding the conventional 0.05 threshold for statistical significance, the moderate effect size (Cohen’s d = 0.626952) indicates a noticeable divergence between the two models. This suggests that with further optimizations, the QLSTM could potentially demonstrate superior performance.
@@
T-Statistic 1.982596
P-Value 0.054677
Effect Size (Cohen’s d) 0.626952
@@
The statistical analysis of the test loss, as depicted in Table 4, unequivocally points to a significant superiority of the QLSTM model. The highly significant p-value (0.000002) accentuates a substantial difference in performance between the two models, with the QLSTM achieving markedly lower test loss values. Furthermore, the pronounced effect size (Cohen’s d = -1.760950) substantiates this claim, portraying the QLSTM as a promising avenue for advancing the state-of-the-art in time series forecasting, specifically in the context of solar power production.
@@
Effect Size (Cohen’s d) -1.760950

### v3

@@
Quantum Long Short-Term Memory (QLSTM) vs Classical LSTM in Time Series Forecasting: A Comparative Study in Solar Power Forecasting
Saad Zafar Khan1 , Nazeefa Muzammil1 , Salman Ghafoor1
Haibat Khan2 , Syed Mohammad Hassan Zaidi3 ,
Abdulah Jeza Aljohani4 , Imran Aziz5
1 SEECS, National University of Sciences and Technology, Islamabad, Pakistan
2 CAE, National University of Sciences and Technology, Islamabad, Pakistan
3 Ghulam Ishaq Khan Institute of Engineering Sciences and Technology, Swabi, Pakistan
4 Department of Electrical and Computer Engineering, King Abdulaziz University, 21589, Jeddah, Saudi Arabia
5 Department of Physics and Astronomy, Uppsala University, 75120 Uppsala, Sweden.
Email: szkhan.bee19seecs@seecs.edu.pkCorresponding author:salman.ghafoor@seecs.edu.pk
Accurate solar power forecasting is pivotal for the global transition towards sustainable energy systems. This study conducts a meticulous comparison between Quantum Long Short-Term Memory (QLSTM) and classical Long Short-Term Memory (LSTM) models for solar power production forecasting. The primary objective is to evaluate the potential advantages of QLSTMs, leveraging their exponential representational capabilities, in capturing the intricate spatiotemporal patterns inherent in renewable energy data. Through controlled experiments on real-world photovoltaic datasets, our findings reveal promising improvements offered by QLSTMs, including accelerated training convergence and substantially reduced test loss within the initial epoch compared to classical LSTMs. These empirical results demonstrate QLSTM’s potential to swiftly assimilate complex time series relationships, enabled by quantum phenomena like superposition. However, realizing QLSTM’s full capabilities necessitates further research into model validation across diverse conditions, systematic hyperparameter optimization, hardware noise resilience, and applications to correlated renewable forecasting problems. With continued progress, quantum machine learning can offer a paradigm shift in renewable energy time series prediction, potentially ushering in an era of unprecedented accuracy and reliability in solar power forecasting worldwide. This pioneering work provides initial evidence substantiating quantum advantages over classical LSTM models while acknowledging present limitations. Through rigorous benchmarking grounded in real-world data, our study illustrates a promising trajectory for quantum learning in renewable forecasting.
Keywords Quantum Machine Learning  $\cdot$ Forecasting  $\cdot$ Quantum Neural Networks  $\cdot$ Renewable energy systems
@@ 1 Introduction
Despite these promising developments, the existing literature significantly lacks a comprehensive, empirical comparison between QLSTM and classical LSTM models grounded in real-world solar production data. Preceding works [1] [7] have primarily focused on synthetic benchmarks or theoretical analysis of Quantum Recurrent Neural Networks variants, leaving a critical gap in our understanding of the practical implications of QLSTMs for renewable energy forecasting. This investigation aims to bridge this divide, offering a thorough comparative analysis in this pivotal domain.
@@ 2 Background or Related Work
The global energy landscape is undergoing a transformative shift towards sustainable and renewable solutions, driven by the pressing need to mitigate the environmental impact of traditional fossil fuel-based systems. Solar power has emerged as a pivotal force in reshaping energy production and consumption patterns, offering a promising pathway to mitigate the environmental impact of traditional fossil fuel-based systems. However, the large-scale integration of intermittent renewable sources into existing infrastructure poses multifaceted challenges that must be addressed to facilitate a seamless and efficient transition. Inaccuracies in forecasting the generation of solar and wind power can lead to significant deviations from planned electricity schedules, resulting in imbalances, inefficiencies, and substantial costs for grid operators and utilities [11]. Consequently, policymakers have prioritized improved renewable forecasting to mitigate such challenges.
@@
Amidst this paradigm shift, the convergence of quantum information (QI) and machine learning (ML) has ushered in a revolutionary approach to data analytics: Quantum Machine Learning (QML) [9]. This paradigm-shifting synthesis harnesses techniques from both quantum computing and traditional machine learning, offering innovative solutions to longstanding obstacles across diverse sectors, including renewable energy [4]. Significantly, QML transcends mere energy minimization tasks, presenting a broader scope in problem-solving paradigms [2], thereby unlocking new avenues for precise solar power predictions.
@@
As the large-scale penetration of renewable sources necessitates proactive management of electrical grids, advanced prediction methodologies for intermittent energy sources, particularly photovoltaic plants, have become paramount [23, 18]. Accurate solar power forecasting plays a critical role in enabling grid operators to maintain a delicate equilibrium between energy creation and utilization. Notably, solar production forecasting over longer time horizons does not mandate real-time predictions, providing an opportunity where the potentially slower inference times of quantum models may be acceptable in exchange for substantially improved accuracy.
@@
In recent years, there has been a surge of interest in quantum machine learning developing and refining of quantum adaptations of recurrent neural networks (RNNs) for time series forecasting applications [19] [15], thereby substantiating the potential of quantum computational models in predictive analytics. Marking a pivotal shift, the seminal work of Chen et al. [8] pioneered the introduction of QLSTM architectures, which amalgamate variational quantum circuits [5] with the conventional LSTM framework. This innovative synthesis harnesses an exponentially larger Hilbert space for data representation and computation, potentially enabling the capture of higher-order correlations and intricate temporal dynamics.
@@
While classical Long Short-Term Memory (LSTM) neural networks have demonstrated remarkable efficacy in leveraging long-term temporal dependencies for accurate forecasting[24] [22] [23] [16], they struggle to capture the complex, non-linear spatiotemporal patterns inherent in solar power generation [14], which involve intricate relationships between meteorological variables, solar irradiance, and power output. Quantum Long Short-Term Memory (QLSTM) networks, which leverage the principles of quantum mechanics and an exponentially larger Hilbert space, hold the potential to address these limitations by enabling more effective mapping of these intricate relationships between weather variables, solar irradiance, and power generation more effectively. and capturing higher-order correlations and temporal dynamics.
@@ 3 Research Motivation and Novelty
This advancement, coupled with the promise of faster convergence times and heightened resilience to noise [10], paves the way for the realization of more precise and reliable forecasting systems – a quality of paramount importance in the domain of solar power production. However, it is crucial to acknowledge that the nascent field of quantum machine learning presents its own unique challenges and opportunities [6]. These challenges encompass the current hardware limitations, the potential trade-offs between quantum advantages and computational overhead, as well as the need for systematic optimization and validation across diverse scenarios, factors that this study meticulously considers in its comparative analysis.
@@ 4 Contributions
  • Empirical Validation on Real-World Data: Providing the first comprehensive empirical evidence that validates the utility of Quantum Long Short-Term Memory (QLSTM) networks for solar power forecasting. This study transitions from synthetic benchmarks to operational photovoltaic plant datasets, offering a realistic evaluation of quantum architectures on genuine renewable time series data. Quantitative results demonstrate that QLSTMs achieve up to 50% improvement in accuracy and 85.7% faster convergence compared to their classical LSTM counterparts.
@@
  • Establishing a Credible Alternative: By benchmarking against classical techniques and conventional neural networks using actual photovoltaic data, this investigation establishes QLSTMs as a credible alternative to traditional forecasting methods. The results accentuate the potential of quantum machine learning in fortifying the accuracy and reliability essential for sustainable energy infrastructure planning and execution, enabling improvements of up to 50% in forecasting accuracy compared to conventional LSTM structure.
@@ 5.1 Data Description
Attribute Real-World Solar Plant Data Simulated Solar Power Data
Source Kaggle [12] NREL’s Solar Power Data for Integration Studies [13]
Geographical Coordinates Operational data from two PV plants in India 33.75° N, 116.65° W (Near Daggett, California)
Capacity Max DC Power: $\sim$298.94 kW Max AC Power: $\sim$29.15 kW 200 MW Utility Scale PV
Duration May 15, 2020 - June 17, 2020 (34 days) Full year of 2006
Resolution 15-minute intervals Power: 5-minute intervals [13] Weather: 30-minute intervals [20]
Attributes Power Output Variables: DC power, AC power, Daily Yield, Total Yield Weather Variables: Ambient temperature, module temperature, irradiation Metadata: Timestamp, plant ID, sensor/inverter ID Power Output Variables: Power (MW) Weather Variables: Temperature, DHI, cloud type, relative humidity, dew point, pressure, windspeed, solar angle Metadata: Datetime
Table 1: Comparison of Real-World and Simulated Solar Power Data
@@ 5.3 Simulation Framework
Our exploration of Quantum Long Short-Term Memory (QLSTM) models was made possible using the PennyLane quantum machine learning framework [3]. PennyLane, at its core, blends quantum and traditional computing to help build and refine models, benefiting from its ability to automatically adjust model parameters.
@@
Bridging the gap between the quantum and regular sections, PennyLane’s qml.qnn.TorchLayer connects the quantum elements with the regular PyTorch framework [17]. This ensures a smooth flow of adjustments during the optimization phase.
@@ 5.4 Architecture
The LSTM and QLSTM architectures are detailed in Appendix. This research encompasses the design and deployment of both LSTM and QLSTM architectures. These architectures were judiciously crafted to allow a fair comparison between classical and quantum techniques, with a focal point on solar power forecasting.
@@
Diverging, the QLSTM model replaces the classical LSTM cells with variational quantum circuits (VQCs), an implementation adapted and enhanced from qlstm repository for parts of speech tagging [21]. This substitution aims to harness the computational advantages unique to quantum mechanisms. Echoing the LSTM’s design, the QLSTM layers two of these quantum circuits. The VQCs, in their capacity as quantum feature extractors, exploit the deep representational capabilities of quantum states, encoding intricate time series dynamics. These parametric circuits oscillate between rotation and entanglement gates, ensuring a concise representation of temporal patterns within the exponentially expansive Hilbert space. The qubit quantity and circuit depth were adapted in alignment with the specific characteristics intrinsic to the solar forecasting data. Notably, outside of its quantum encoding, the QLSTM’s broader workflow aligns seamlessly with the LSTM’s, facilitating a direct comparison. A dropout rate of 0.2 is also infused between the QLSTM layers, maintaining consistency.
@@
  • Quantum Variational Circuit (VQC): Data is encoded into the quantum circuit using rotation operations, followed by alternating entanglement and variational rotation layers detailed in Appendix.
@@
1 Input layer Input layer
2 LSTM layer (with input, output, forget gates and memory cell) QLSTM layer (with VQCs featuring rotation and entanglement gates)
3 Dropout (0.2 rate) Dropout (0.2 rate)
4 LSTM layer (with input, output, forget gates and memory cell) QLSTM layer (with VQCs featuring rotation and entanglement gates)
5 Dropout (0.2 rate) Dropout (0.2 rate)
6 Linear output layer Linear output layer
Table 2: Comparison of LSTM and QLSTM Architectures
5.5 Encoding Process
The process of encoding classical data into quantum states is a crucial component of our QLSTM model. This section details the steps involved in this encoding process and analyzes its computational complexity.
5.5.1 Encoding Steps
The encoding process consists of three main steps:
Data Preprocessing
Before quantum encoding, we preprocess the input data to obtain suitable rotation angles for quantum gates:
  • ry_params: Computed as $\arctan(\text{feature})$ for each input feature.
  • rz_params: Computed as $\arctan(\text{feature}^{2})$ for each input feature.
Quantum State Preparation
For each qubit in the circuit, we apply the following gates sequentially:
  1. 1.
    Hadamard gate (qml.Hadamard): Creates an initial superposition state.
  2. 2.
    RY rotation (qml.RY): Applied using the preprocessed ry_params.
  3. 3.
    RZ rotation (qml.RZ): Applied using the preprocessed rz_params.
Variational Quantum Circuit
After state preparation, we apply a variational quantum circuit. This circuit, implemented through the ansatz function, is repeated n_qlayers times and includes:
  • Entangling operations: CNOT gates between qubits.
  • Additional rotations: RX, RY, and RZ gates with learnable parameters.
5.5.2 Complexity Analysis
We analyze the complexity of our encoding process in terms of both classical preprocessing and quantum operations.
Classical Preprocessing Complexity
  • Computing ry_params and rz_params: $\mathcal{O}(n_{\text{features}})$
This step scales linearly with the number of input features.
Quantum State Preparation Complexity
  • Applying Hadamard, RY, and RZ gates: $\mathcal{O}(n_{\text{qubits}})$
This step scales linearly with the number of qubits.
Variational Quantum Circuit Complexity
  • Entangling layer: $\mathcal{O}(n_{\text{qubits}}^{2}\cdot n_{\text{qlayers}})$
    • Each layer applies $\mathcal{O}(n_{\text{qubits}})$ CNOT gates, repeated for each qubit.
  • Variational layer: $\mathcal{O}(n_{\text{qubits}}\cdot n_{\text{vrotations}}\cdot n_{\text{qlayers}})$
    • Each qubit undergoes n_vrotations rotations in each of the n_qlayers.
Overall Encoding Complexity
The total complexity of encoding data into quantum states is:
$\mathcal{O}(n_{\text{features}}+n_{\text{qubits}}+n_{\text{qubits}}^{2}\cdot n_{\text{qlayers}}+n_{\text{qubits}}\cdot n_{\text{vrotations}}\cdot n_{\text{qlayers}})$
In most practical scenarios, this can be simplified to:
$\mathcal{O}(n_{\text{qubits}}^{2}\cdot n_{\text{qlayers}})$
as the number of qubits and layers typically dominates the complexity.
@@ 6 Results
The metrics utilized for the statistical and predictive analysis, are defined in detail in the Appendix (Evaluation Methodology).
@@ 6.1 Statistical Analysis
In this section, we conducted a statistical analysis to compare the losses between the QLSTM model and the classical LSTM model as shown in Table 5
@@
The statistical analysis of the train loss, as shown in Table 3, suggests that the QLSTM model tends to perform better in terms of reducing train loss, as in table . Although the p-value (0.0547) is slightly above the conventional threshold for statistical significance (0.05), the moderate effect size (Cohen’s d = 0.627) indicates a noticeable difference favoring the QLSTM model. This suggests that with further refinement, the QLSTM could consistently outperform the classical LSTM in training performance.
@@
T-Statistic 1.983
P-Value 0.0547
Effect Size (Cohen’s d) 0.627
@@
Similarly, we compared the test losses between the QLSTM and classical LSTM models to assess how well each model generalizes to unseen data, as depicted in Table 4. The results, as presented in Table 4, clearly show that the QLSTM model significantly outperforms the classical LSTM in terms of test loss. The highly significant p-value (0.000002) and the large effect size (Cohen’s d = -1.761) indicate that the QLSTM model achieves much lower test losses, making it a more effective model for generalization and prediction accuracy in time series forecasting, particularly for solar power production.
@@
Effect Size (Cohen’s d) -1.761